Barn painting. Melanie can Saint a certain barn by herself in x days. Her helper Melissa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x = 5
Answer by stanbon(75887)

Barn painting. Melanie can Saint a certain barn by herself in x days. Her helper Melissa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x = 5
------------
Melanie DATA:
Time = x days/job ; Rate = 1/x job/day
-----------
Melissa DATA:
Time = 2x days/job ; Rate = 1/2x job/day
-----------------
Together data:
Rate= 1/x + 1/2x = 2/2x+1/2x= 3/2x job/day
---------
If x=5, Rate = 3/(2*5)=3/10 job in a day
Cheers,
Stan H.